# An Extension of the Localized Artificial Diffusivity Method for Immiscible and High Density Ratio Flows

### v1

@@ Abstract
The localized artificial diffusivity (LAD) method is widely regarded as the preferred multi-material regularization scheme for the compact finite difference method, because it is conservative, easy to implement, and generally robust for a wide range of multi-material problems. However, traditional LAD methods face significant challenges when applied to flows with large density ratios and when maintaining thermodynamic equilibrium across material interfaces. These limitations arise from the formulation of the artificial diffusivity flux and the reliance on enthalpy diffusion for interface regularization. Additionally, traditional LAD methods struggle to ensure stability under large density ratio conditions, fail to maintain a finite interface thickness, and are therefore unsuitable for modeling immiscible interfaces. In this work, we discuss the origins of these issues in traditional LAD methods and propose modifications which enable the simulation of large density ratio and immiscible flows. The proposed method targets the artificial diffusion fluxes at gradients and ringing in the volume fraction, rather than the mass fraction in traditional methods, to consistently regularize large density ratio interfaces. Furthermore, the proposed method introduces an artificial bulk density diffusion term to enforce equilibrium conditions across interfaces. To address the challenge of modeling immiscible flows, a conservative diffuse interface term is incorporated into the formulation to ensure a finite interface thickness. Specific consideration is taken in the design of the method to ensure that these crucial properties are maintained for $N$-material flows. The effectiveness of the proposed method is demonstrated through a series of canonical test cases, and its accuracy is validated by comparison with experimental data on microbubble collapse in water. These results highlight the method’s robustness and its ability to overcome the limitations of traditional LAD approaches.
@@ 1 Introduction
Simulations of multi-material flows are often performed on Eulerian grids, Lagrangian frames, or with arbitrary Lagrangian-Eulerian (ALE) methods. Lagrangian and ALE methods have shown successes for capturing complex multi-material mixing, but tend to be diffusive and suffer from poorly conditioned cells and mesh tangling [6, 7]. In this work, we use Eulerian grids because they allow for the use of high-order finite difference methods, which can capture the small-scale features of turbulence without over dissipation [8]. In the context of Eulerian methods, there are two main multi-material approaches: interface tracking and interface detection. Interface tracking involves explicitly evolving the location of the interface in methods such as volume-of-fluid [9] and level-set methods [10]. These methods accurately capture interface shapes but require special treatments at the interfaces and struggle to maintain exact conservation [11]. Interface detection methods compute the location of the interface using the flow variables within the simulation, which allows for better capturing of topology changes and interface creation. Within interface detection methods, there are sharp and diffuse interface methods. Sharp interface methods capture the material interface as a single-point jump, which often requires special treatment to maintain numerical stability [12]. Diffuse interface methods spread the material interface over multiple grid points, which avoids numerical instabilities, but introduces complexity with the treatment of points with multiple materials [13, 5, 14, 15]. In this work, we are using a diffuse interface approach due to its simplicity of interface representation and conservation properties.
@@
Another weakness of LAD methods is that they continuously diffuse at the interface, so a finite interface thickness cannot be enforced. One common way to address this is through the phase-field method, where an additional PDE is solved to compute an additional flux term to enforce constant a constant interface thickness [31]. This method has seen success for a wide range of multi-material problems [29, 32, 33, 18, 34, 35, 36, 37]. A particular class of phase-field methods is the conservative diffuse interface (CDI) methods, where a conservative flux term is used to force multi-material interfaces toward a finite thickness while maintaining conservation [38, 16]. These methods have been successful at ensuring conservation while maintaining fixed interface thicknesses for immiscible interfaces withing a variety of numerical methods [38, 39, 40, 33, 41]. The CDI method has also been applied in the context of four-equation models where the volume fraction is computed from the conserved variables and is used as the phase-field variable without the addition of an extra PDE [25, 42].
@@ 2 Governing Equations
where $t$ is time, $\rho=\sum_{i=1}^{N_{s}}V_{i}\rho_{i}$ is the mixture density, $N_{s}$ is the number of species, $V_{i}$ is the volume fraction of material $i$, $\rho_{i}$ is the density of material $i$, $Y_{i}$ is the mass fraction of material $i$, $\bm{u}$ is the velocity, $J_{i}$ is the species diffusion flux, $p$ is pressure, $\bm{\tau}$ is the stress tensor, $\bm{F}$ is the momentum consistency flux, $E$ is total energy, $\bm{q}$ is the thermal conductivity flux, $\bm{q}_{d}$ is the enthalpy flux between materials, and $\bm{H}$ is the energy consistency flux. The total energy is defined as
@@
The artificial fluxes will be defined in Section 3. The materials each have their own equation of state. In mixed grid points, the materials are assumed to the in pressure-temperature equilibrium, which is found using a Newton-iteration within each point as describe in [19].
@@ 3 Numerical Method
[…]
@@ 3.1 Bubble Advection Stability Test
In order to study and demonstrate the limitations of the traditional multi-material methods for large density ratios, we introduce this test case. In this case, a high density ratio bubble is advected through a domain of two other materials. The nondimensional 2D domain is $[0,1]\times[0,1]$ and periodic in both directions. A circular bubble with radius $r=0.25$ of material $h$ with density $\rho_{h}=R$ is initialized in the center of the domain. Outside of the bubble, the domain is split into two sections. For $y>0.5$, material $l$ is defined with density $\rho_{l}=1$. For $y<0.5$, material $m$ is defined with density $\rho_{m}=10^{\log_{10}(R/2)}$. All three materials are $\gamma$-law gases with $\gamma=1.4$ and $c_{v,i}=\frac{1}{\rho_{i}(\gamma-1)}$. The interfaces between the materials are defined smoothly using a hyperbolic tangent
@@
where $N_{p}$ is a chosen number of grid points in the interface thickness and $\Delta$ is the grid spacing. This form is chosen so that 99% of the interface is contained within $N_{p}$ grid points centered around $r_{0}$. The number of grid points in the domain varies with $N_{p}$ such that the physical thickness of the interface is $0.05$ to prevent the two sides of the bubble from interacting. An image of the initial domain is show in Fig 1. The entire domain is initialized at equilibrium with $P=1$, $T=1$, $\bm{u}=[10,0]$. The Mach number in the heavy material is $M=u_{1}/\sqrt{\gamma P/R}=10\sqrt{R/\gamma}$, which gives a range of $M=26.73-8451$ for the density ratios of $R=10-10^{6}$ that are tested. The bubble is advected for $10$ periods. Whether the simulation ran to completion and the magnitude of the pressure oscillations at completion are reported. The study was also repeated with larger artificial diffusivities and lower timesteps. These changes show some improvement for the stability of the method for larger density ratios, but do not affect the overall trends. Hence, the study using only the default coefficients is reported.
@@ 3.2 Limitations of LAD Formulation
Additionally, for large density ratios, the method fails to regularize the interface and often goes unstable. To demonstrate this, we will study the advection of a high density bubble as detailed in Sec. 3.1. Figure 2 shows the results of this stability test for a range of $R$ and $N_{p}$. Empty markers indicate that the simulation went unstable before reaching completion. We see that for lower density ratios, the method remains stable for even $N_{p}=8$, while for $R>1000$, increasing numbers of points are needed to remain stable. Also, the pressure oscillations are at best around 1% of the pressure and at worst greater than 10%. This is because the method does not maintain equilibrium, as previously discussed. This performance leaves much to be desired because thick interfaces are necessary to stabilize large density ratios and the pressure oscillations are quite large, even in the best cases.
@@
4 Localized Artifical Diffusivity Method
@@
$\bm{H}$ $=\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right)\left(\frac{1}{2}\rho\bm{u}\cdot\bm{u}\right),$ (39)
@@
A single diffusivity is used in all of the materials in order to prevent pressure oscillations, which will be explained in detail in Sec. 4.1. In cases with two materials, the balance between the two mass fractions naturally created equal diffusivities. However, in the $N_{s}>2$ case, it is necessary to force the diffusivities to be equal. The overshoot diffusivities, $D^{*}_{D,Y}$ and $D^{*}_{D,V}$, are combined using the maximum so that the diffusivity addresses the worst overshoots. The ringing diffusivities, $D^{*}_{O,Y}$ and $D^{*}_{O,V}$, are volume fraction weighted to avoid large amounts of diffusivity being added when only trace amounts of a material are present.
@@ 4.1 Equilibrium
$\frac{\partial\rho Y_{i}}{\partial t}$ $+u_{0}\nabla(\rho Y_{i})=-\nabla\bm{J}_{i},$ (41)
$\frac{\partial\rho u}{\partial t}$ $+u_{0}^{2}\nabla\rho=-u_{0}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right),$ (42)
$\frac{\partial E}{\partial t}$ $+u_{0}\nabla E=-\nabla\cdot\bm{q}_{d}-\frac{1}{2}u_{0}^{2}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right).$ (43)
@@
$\frac{\partial\rho}{\partial t}$ $=-u_{0}\nabla\rho-\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right),$ (44)
$\frac{\partial\rho Y_{i}}{\partial t}$ $=-u_{0}\nabla(\rho Y_{i})-\nabla\bm{J}_{i},$ (45)
$\frac{\partial\rho u}{\partial t}$ $=-u_{0}^{2}\nabla\rho-u_{0}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right),$ (46)
$\frac{\partial E}{\partial t}$ $=-\frac{1}{2}u_{0}^{3}\nabla\rho-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}-\frac{1}{2}u_{0}^{2}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right).$ (47)
@@
$=\frac{1}{\rho}\left(-u_{0}^{2}\nabla\rho-u_{0}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right)\right)-\frac{u_{0}}{\rho}\left(-u_{0}\nabla\rho-\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right)\right),$ (53)
$=\left(\frac{u_{0}^{2}}{\rho}-\frac{u_{0}^{2}}{\rho}\right)\nabla\rho+\left(\frac{u_{0}}{\rho}-\frac{u_{0}}{\rho}\right)\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right),$ (54)
@@
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\frac{1}{2}u_{0}^{3}\nabla\rho-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}-\frac{1}{2}u_{0}^{2}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right)-\frac{1}{2}u_{0}^{2}\left(-u_{0}\nabla\rho-\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}\right)\right)-\sum_{i=1}^{N_{s}}e_{i}\frac{\partial\rho Y_{i}}{\partial t}\right],$ (58)
@@
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}-\sum_{i=1}^{N_{s}}e_{i}\left(-u_{0}\nabla(\rho Y_{i})-\nabla\bm{J}_{i}\right)\right],$ (60)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}+u_{0}\nabla\left(\sum_{i=1}^{N_{s}}\rho Y_{i}e_{i}\right)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}e_{i}\right)\right],$ (61)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}+u_{0}\nabla(\rho e)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}e_{i}\right)\right],$ (62)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\bm{q}_{d}+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}e_{i}\right)\right],$ (63)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}e_{i}\right)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}e_{i}\right)\right],$ (64)
@@
$\frac{\partial p}{\partial t}$ $=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}h_{i}\right)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}_{i}e_{i}\right)\right],$ (66)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}i\frac{P_{0}}{\rho_{i}}\right)\right],$ (67)

### v2

@@ Abstract
The localized artificial diffusivity (LAD) method is widely regarded as the preferred multi-material regularization scheme for the compact finite difference method, because it is conservative, easy to implement, and generally robust for a wide range of multi-material problems. However, traditional LAD methods face significant challenges when applied to flows with large density ratios and when maintaining thermodynamic equilibrium across material interfaces. These limitations arise from the formulation of the artificial diffusivity flux and the reliance on enthalpy diffusion for interface regularization. Additionally, traditional LAD methods struggle to ensure stability under large density ratio conditions, fail to maintain a finite interface thickness, and are therefore unsuitable for modeling immiscible interfaces. In this work, we discuss the origins of these issues in traditional LAD methods and propose modifications which enable the simulation of large density ratio and immiscible flows. The proposed method targets the artificial diffusion fluxes at gradients and ringing in the volume fraction, rather than the mass fraction in traditional methods, to consistently regularize large density ratio interfaces. Furthermore, the proposed method introduces an artificial bulk density diffusion term to enforce equilibrium conditions across interfaces. To address the challenge of modeling immiscible flows, a conservative diffuse interface term is incorporated into the formulation to ensure a finite interface thickness. Specific consideration is taken in the design of the method to ensure that these crucial properties are maintained for $N$-material flows. The effectiveness of the proposed method is demonstrated through a series of canonical test cases, and its accuracy is validated by comparison with experimental data on micro-bubble collapse in water. These results highlight the method’s robustness and its ability to overcome the limitations of traditional LAD approaches.
@@ 1 Introduction
Simulations of multi-material flows are often performed on Eulerian grids, Lagrangian frames, or with arbitrary Lagrangian-Eulerian (ALE) methods. Lagrangian and ALE methods have shown successes for capturing complex multi-material mixing, but tend to be diffusive and suffer from poorly conditioned cells and mesh tangling [6, 7]. In this work, we use Eulerian grids because they allow for the use of high-order finite difference methods, which can capture the small-scale features of turbulence without over-dissipation [8]. In the context of Eulerian methods, there are two main multi-material approaches: interface tracking and interface detection. Interface tracking involves explicitly evolving the location of the interface in methods such as the marker-in-cell method [9] and front-tracking methods [10]. These methods accurately capture interface shapes, but require complex treatments at the interface. Interface detection methods compute the location of the interface using the flow variables within the simulation, which allows for better capturing of topology changes and interface creation. Within interface detection methods, there are sharp and diffuse interface methods. Sharp interface methods capture the material interface as a single-point jump, which often requires special treatment to maintain numerical stability [11]. Additionally, methods such as the volume-of-fluid [12] and level-set methods [13] reconstruct the interface location for accuracy, but struggle to maintain exact conservation [14]. Diffuse interface methods spread the material interface over multiple grid points, which avoids numerical instabilities, but introduces complexity with the treatment of points with multiple materials [15, 5, 16, 17]. In this work, we are using a diffuse interface approach due to its simplicity of interface representation and conservation properties.
@@
Another weakness of LAD methods is that they continuously diffuse at the interface, so a finite interface thickness cannot be enforced. One common way to address this is through the phase-field method, where an additional PDE is solved to compute an additional flux term to enforce a constant interface thickness [33]. This method has seen success for a wide range of multi-material problems [31, 34, 35, 20, 36, 37, 38, 39]. The conservative diffuse interface (CDI) method is a particular choice of phase-field model where a conservative flux term is used to force multi-material interfaces toward a finite thickenss while maintaining conservation [40, 41, 18]. The accurate conservative diffuse interface (ACDI) method is another conservative phase-field model that formulates its sharpening flux with a signed distance function for improved accuracy [42]. These methods have been successful at ensuring conservation while maintaining fixed interface thicknesses for immiscible interfaces within a variety of numerical methods [41, 42, 43, 35, 44, 39, 45]. The CDI and ACDI methods have also been applied in the context of four-equation models where the volume fraction is computed from the conserved variables and is used as the phase-field variable without the addition of an extra PDE [27, 46, 47].
@@ 2 Governing Equations
where $t$ is time, $\rho=\sum_{i=1}^{N_{s}}V_{i}\rho_{i}$ is the mixture density, $N_{s}$ is the number of species, $V_{i}$ is the volume fraction of material $i$, $\rho_{i}$ is the density of material $i$, $Y_{i}$ is the mass fraction of material $i$, $\bm{u}$ is the velocity, $\bm{J}_{i}$ is the species diffusion flux, $p$ is pressure, $\bm{\tau}$ is the stress tensor, $\bm{F}$ is the momentum consistency flux, $E$ is total energy, $\bm{q}$ is the thermal conductivity flux, $\bm{q}_{d}$ is the enthalpy flux between materials, and $\bm{H}$ is the energy consistency flux. The total energy is defined as
@@
The artificial fluxes will be defined in Section 3. The separation of physical and numerical viscous fluxes allows the formulation to capture a wide range of scenarios. For example, in multi-material simulations of materials that diffuse into each other, $D_{f,i}$ is nonzero, but when that is not the case, $D_{f,i}$ is set to zero and only the artificial diffusion is used for stabilization. The materials each have their own equation of state. In mixed grid points, the materials are assumed to the in pressure-temperature equilibrium, which is found using a Newton-iteration within each point as described in [21].
@@ 3 Numerical Method
$G(\zeta;L)$ $=\frac{e^{-6\zeta^{2}/L^{2}}}{\int_{-L}^{L}e^{-6\zeta^{2}/L^{2}}d\zeta},\quad L=4\Delta.$ (29)
@@ 3.1 Bubble Advection Stability Test
In order to study and demonstrate the limitations of the traditional multi-material methods for large density ratios, we introduce this test case. In this case, a high density ratio bubble is advected through a domain of two other materials. The nondimensional 2D domain is $0,1]\times[0,1]$ and periodic in both directions. A circular bubble with radius $r=0.25$ of material $h$ with density $\rho_{h}=R$ is initialized in the center of the domain. Outside of the bubble, the domain is split into two sections. For $y>0.5$, material $l$ is defined with density $\rho_{l}=1$. For $y<0.5$, material $m$ is defined with density $\rho_{m}=10^{\log_{10}(R)/2}$. All three materials are $\gamma$-law gases with $\gamma=1.4$ and $c_{v,i}=\frac{1}{\rho_{i}(\gamma-1)}$. The interfaces between the materials are defined smoothly using a hyperbolic tangent
@@
where $N_{p}$ is a chosen number of grid points in the interface thickness and $\Delta$ is the grid spacing. This form is chosen so that 99% of the interface is contained within $N_{p}$ grid points centered around $r_{0}$. The number of grid points in the domain varies with $N_{p}$ such that the physical thickness of the interface is $0.05$ to prevent the two sides of the bubble from interacting. An image of the initial domain is shown in Fig 1. The entire domain is initialized at equilibrium with $P=1$, $T=1$, $\bm{u}=[10,0]$. The Mach number in the heavy material is $M=u_{1}/\sqrt{\gamma P/R}=10\sqrt{R/\gamma}$, which gives a range of M=26.738451M=26.73-8451 for the density ratios of $R=10-10^{6}$ that are tested. The bubble is advected for $10$ periods. Whether the simulation ran to completion and the magnitude of the pressure oscillations at completion are reported. Simulations that did not run to completion fail because ringing at the material interface results in an imaginary speed of sound. The study was also repeated with larger artificial diffusivities and lower timesteps. These changes show some improvement for the stability of the method for larger density ratios, but do not affect the overall trends. Hence, the study uses the default LAD coefficients from Sec. 3 and a CFL of 0.2 is used in order keep the timesteps comparable between different methods.
@@ 3.2 Limitations of LAD Formulation
Additionally, for large density ratios, the method fails to regularize the interface and often goes unstable. To demonstrate this, we will study the advection of a high density bubble as detailed in Sec. 3.1. Figure 2 shows the results of this stability test for a range of $R$ and $N_{p}$. Empty markers indicate that the simulation went unstable before reaching completion. We see that for lower density ratios, the method remains stable for even $N_{p}=7$, while for $R>1000$, increasing numbers of points are needed to remain stable. Also, the pressure oscillations are at best around 1% of the pressure and at worst greater than 25%. This is because the method does not maintain equilibrium, as previously discussed. This performance leaves much to be desired because thick interfaces are necessary to stabilize large density ratios and the pressure oscillations are quite large, even in the best cases.
@@
4 Localized Artificial Diffusivity Method
@@
$\bm{H}$ $=\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right)\left(\frac{1}{2}\bm{u}\cdot\bm{u}\right),$ (39)
@@
A single diffusivity is used in all of the materials in order to prevent unphysical oscillations, which will be explained in detail in Sec. 4.3. In cases with two materials, the balance between the two mass fractions naturally created equal diffusivities. However, in the $N_{s}>2$ case, it is necessary to force the diffusivities to be equal. The overshoot diffusivities, $D^{*}_{D,Y}$ and $D^{*}_{D,V}$, are combined using the maximum so that the diffusivity addresses the worst overshoots. The ringing diffusivities, $D^{*}_{O,Y}$ and $D^{*}_{O,V}$, are volume fraction weighted to avoid large amounts of diffusivity being added when only trace amounts of a material are present.
@@ 4.1 Equilibrium
$\frac{\partial\rho Y_{i}}{\partial t}$ $+u_{0}\nabla(\rho Y_{i})=-\nabla\bm{J}^{*}_{i},$ (41)
$\frac{\partial\rho u}{\partial t}$ $+u_{0}^{2}\nabla\rho=-u_{0}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right),$ (42)
$\frac{\partial E}{\partial t}$ $+u_{0}\nabla E=-\nabla\cdot\bm{q}_{d}-\frac{1}{2}u_{0}^{2}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right).$ (43)
@@
$\frac{\partial\rho}{\partial t}$ $=-u_{0}\nabla\rho-\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right),$ (44)
$\frac{\partial\rho Y_{i}}{\partial t}$ $=-u_{0}\nabla(\rho Y_{i})-\nabla\bm{J}^{*}_{i},$ (45)
$\frac{\partial\rho u}{\partial t}$ $=-u_{0}^{2}\nabla\rho-u_{0}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right),$ (46)
$\frac{\partial E}{\partial t}$ $=-\frac{1}{2}u_{0}^{3}\nabla\rho-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}-\frac{1}{2}u_{0}^{2}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right).$ (47)
@@
$=\frac{1}{\rho}\left(-u_{0}^{2}\nabla\rho-u_{0}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}\right)\right)-\frac{u_{0}}{\rho}\left(-u_{0}\nabla\rho-\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right)\right),$ (53)
$=\left(\frac{u_{0}^{2}}{\rho}-\frac{u_{0}^{2}}{\rho}\right)\nabla\rho+\left(\frac{u_{0}}{\rho}-\frac{u_{0}}{\rho}\right)\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right),$ (54)
@@
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\frac{1}{2}u_{0}^{3}\nabla\rho-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}-\frac{1}{2}u_{0}^{2}\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right)-\frac{1}{2}u_{0}^{2}\left(-u_{0}\nabla\rho-\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right)\right)-\sum_{i=1}^{N_{s}}e_{i}\frac{\partial\rho Y_{i}}{\partial t}\right],$ (58)
@@
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}-\sum_{i=1}^{N_{s}}e_{i}\left(-u_{0}\nabla(\rho Y_{i})-\nabla\bm{J}^{*}_{i}\right)\right],$ (60)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}+u_{0}\nabla\left(\sum_{i=1}^{N_{s}}\rho Y_{i}e_{i}\right)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}e_{i}\right)\right],$ (61)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-u_{0}\nabla(\rho e)-\nabla\cdot\bm{q}_{d}+u_{0}\nabla(\rho e)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}e_{i}\right)\right],$ (62)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\bm{q}_{d}+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}e_{i}\right)\right],$ (63)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}e_{i}\right)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}e_{i}\right)\right],$ (64)
@@
$\frac{\partial p}{\partial t}$ $=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}h_{i}\right)+\nabla\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}e_{i}\right)\right],$ (66)
$=\left(\frac{\partial\rho e}{\partial p}\right)_{\rho Y_{i}}^{-1}\left[-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\frac{P_{0}}{\rho_{i}}\right)\right],$ (67)